Solve using the five-step problem-solving process. Show all neccessary steps to arrive at your solution. Round your answer to 2 decimal places.
A cylindrical drum has a diameter of 3 yards and a volume of 65 cubic yards. What is the height of the drum? (Hint: Use V = ∏r^2h)
I have no idea where to start.Please help me.

Answer by edjones(8007)

pi*r^2h=V
pi*1.5^2*h=65
pi*2.25*h=65
h=65/(pi*2.25)
h=65/7.07
h=9.19 yds
.
Ed